OpenStudy (anonymous):

Can someone please check these 3 Algebra1 problems with their answer please and tell if they are right? x + 12 over x + 4 = x over x + 8 is the answer x = 6 x^2 -1 minus x - 1 ____________minus ________ x^2 - x - 2 x-2 is the answer -1 or 1 x^2 + 9x + 20 divided x + 4 _______________divided________ x^2 - 25 x-4 is the answer x-4 over x-5 Thank you!!!

OpenStudy (kinggeorge):

Let's go through these one at a time. Your first answer is almost correct. Instead of 6, it should be -6.

OpenStudy (anonymous):

OH okay I thought so

OpenStudy (kinggeorge):

For the second one, is it\[{x^2-1 \over x^2-x-2} - {x-1 \over x-2}\]Or\[{x^2-1 \over x^2-x-2} = {x-1 \over x-2}\]

OpenStudy (anonymous):

first one (:

OpenStudy (kinggeorge):

Alright, long story short, I'm not getting \(\pm1\) for an answer.

OpenStudy (anonymous):

oh okay so it must be 0 or x-1 over x-2

OpenStudy (kinggeorge):

First step would be to multiply \[{x-1 \over x-2} \cdot {x+1 \over x+1}\]On the Right side of your equation.

OpenStudy (kinggeorge):

So you end up getting \[{x^2-1 \over x^2-x-2} - {x^2-1 \over x^2-x-2}=0\]So your second answer was correct.

OpenStudy (anonymous):

oh okay zero is correct?

OpenStudy (kinggeorge):

For the last one then it's \[\Large { {{x^2 + 9x + 20} \over {x^2-25}} \over {{x + 4} \over x-4}}\]Correct? And 0 is correct for the previous problem.

OpenStudy (anonymous):

yes !!

OpenStudy (kinggeorge):

You solution for that is correct as well. It should be \({(x-4) / (x-5)}\)

OpenStudy (anonymous):

okay Thank you so much @KingGeorge !!!!!

OpenStudy (kinggeorge):

You're welcome.